# The Class 0 protostars in Orion: Characterizing the properties of their magnetized envelopes

### v1

@@ 3.1 Physical Envelope Properties
HH212M 14.00 0.52 29.1 18.2 0.32 1.6E+23 6.6E+06 0.73 / 1.00 103 78.6 /
HOPS-10 3.33 0.17 21.8 16.5 0.12 6.2E+22 2.6E+06 0.68 / 1.00 53 46.8 /
HOPS-12E 7.31 0.33 24.8 16.5 0.21 1.0E+23 4.3E+06 0.78 / 2.50 62 58.7 /
HOPS-50 4.20 0.32 22.4 18.8 0.23 1.2E+23 4.8E+06 1.01 / 1.00 173 82.8 /
HOPS-53 26.42 0.37 34.1 21.2 0.15 7.6E+22 3.2E+06 0.74 / 1.00 33 21.6 /
HOPS-60 21.93 0.55 32.6 22.9 0.24 1.2E+23 5.0E+06 0.87 / 1.00 153 62.7 /
HOPS-81 1.24 0.14 20.5 17.9 0.11 5.6E+22 2.4E+06 0.74 / 1.00 66 32.4 /
HOPS-84 49.11 0.59 39.8 20.5 0.20 1.1E+23 4.2E+06 0.92 / 2.00 159 87.4 /
HOPS-153 4.43 0.22 22.5 14.0 0.16 7.8E+22 3.3E+06 0.92 / 1.00 103 129.1 /
HOPS-203N 20.44 0.42 32.0 15.7 0.18 9.0E+22 3.8E+06 0.86 / 3.00 31 58.6 /
HOPS-203S 20.44 0.38 32.0 15.7 0.16 8.1E+22 3.4E+06 0.71 / 3.00 33 30.9 /
HOPS-247 3.09 0.37 21.6 13.1 0.34 1.7E+23 7.1E+06 0.84 / 1.00 227 85.3 /
HOPS-340 1.85 0.21 20.8 14.7 0.20 1.0E+23 4.2E+06 0.70 / 1.00 60 41.7 /
HOPS-354 6.57 0.34 24.2 / 0.18 9.1E+22 3.8E+06 0.90 / 1.00 30 24.4 /
HOPS-358 24.96 0.44 33.7 17.6 0.22 1.1E+23 4.6E+06 1.23 / 1.00 111 136.3 /
HOPS-383 7.83 0.22 25.2 20.3 0.14 6.9E+22 2.9E+06 0.89 / 1.00 96 61.7 /
Column 2 to 14 present the bolometric luminosity ($L_{\rm{bol}}$), total continuum flux density at 345 GHz ($S_{\nu}$), averaged envelope dust temperature ($\overline{T}_{\rm env}$), cloud dust temperature ($T_{\rm c}$), envelope gas mass ($M_{\rm env}$), gas column density ($N_{\rm gas}$), gas number density ($n_{\rm gas}$), non-thermal velocity dispersion ($\sigma_{\rm nth}$), B-field angle dispersion ($\delta\phi$), fragmentation level ($N_{F}$), disk mass ($M_{\rm disk}$), disk radius ($R_{\rm disk}$), and the misalignment between outflow and B-field ($\Delta\theta=|\theta_{O}-\theta_{B}|$), respectively. “Standard”, “Rotated”, “Spiral”, “Complex” indicate protostars with different types of B-field structure, “Unresolved” means the B-field pattern is not sufficient to identify its morphology, while “Scattering” represents protostars with compact polarized emission, which is thought to be produced by self-scattering (Huang et al., 2024).$\dagger$\$$\dagger$\$footnotemark: Typical uncertainties for the different parameters are: 10% for $S_{\nu}$; 20–30% for $\overline{T}_{\rm env}$; 30–40% for $M_{\rm env}$, $N_{\rm gas}$, and $n_{\rm gas}$; 0.37 km s-1 for $\sigma_{\rm nth}$.aafootnotemark: The disk emission is contaminated by the envelope dust emission.bbfootnotemark: The disk mass, radius, and their uncertainties are obtained from Tobin et al. (2020).ccfootnotemark: The misalignment between outflow and B-field and its uncertainty are obtained from Huang et al. (2024).
@@
In the case of a dusty cloud heated by internal stellar radiation, the radial temperature profile can be expressed as a power law: $T_{\rm dust}\propto r^{-2/(4+\beta)}$ (Kenyon et al., 1993), where $\beta$ is the dust opacity spectral index. Girart et al. (2009) adopted a value of $\beta=1.5$ when studying low-mass protostars at 1000 au scales. While Kwon et al. (2009) suggested that $\beta$ should be around or smaller than 1 for Class 0 stage. Tobin et al. (2020) fitted a temperature power-law index of -0.46 for hundreds-au-size disks within the Orion star-forming protostellar system, corresponding to $\beta\sim 0.35$, scaled by luminosity. However, this profile shows a slight decline in the outer 100 au regions due to backheating from the surrounding envelope (Tobin et al., 2020), it is subsequently recommended to refine the index to -0.50 (Sheehan et al., 2022). For this study, we use the temperature model proposed by Tobin et al. (2020) for Orion star-forming regions, with $\beta$ set to 1 (corresponding to the index of $\sim-0.40$) for the Class 0 protostellar envelope with a radius of 1200 au. Therefore, the dust temperature $T_{\rm dust}$ can be expressed as:
@@
where $T_{0}=43$ K is the average temperature for a $\sim$ 1 $L_{\odot}$ protostar at a radius of $\sim$50 au (Whitney et al., 2003; Tobin et al., 2013, 2020), $r$ is the distance from the protostar, and $L_{\rm bol}$ is the bolometric luminosity taken from Tobin et al. (2020) and Federman et al. (2023). Given the clear detection of C17O emission in our sample, it is inferred that the dust temperature cannot fall below 20 K, as CO would otherwise be frozen onto the dust grains, making it undetectable (e.g., Jørgensen et al., 2015). The Equation 2 may yield very low dust temperatures (below 20 K) at the edge of the selected regions, especially when the bolometric luminosity $L_{\rm bol}$ is very low. In such cases, we use 20 K as the minimum dust temperature for pixels where the calculated temperature falls below this value. Excluding the extremely high temperature in the central pixel, we calculate the average envelope dust temperature $\overline{T}_{\rm env}$ within inner regions with a radius of $R=1200$ au.
Lombardi et al. (2014) used Planck and Herschel observations of Orion to construct and fit SEDs, and further to obtain a map of dust temperature. As shown in Figure 1, comparing the cloud dust temperature $T_{\rm c}$ (Lombardi et al., 2014) with the envelope dust temperature $\overline{T}_{\rm env}$ (this work, derived using our method described above), we find that 3 protostars have values of $\overline{T}_{\rm env}$ comparable to the dust temperature of the surrounding molecular cloud (near the grey dashed line in Figure 1). Among these, two protostars (HOPS-325 and HOPS-409) exhibit a rotated-hourglass structure of the B-field, while one (HOPS-81) shows compact polarized emission (self-scattering). The remaining protostars have $\overline{T}_{\rm env}$ values higher than $T_{\rm c}$ (above the grey dashed line in Figure 1). Given that the scales of molecular clouds in question are on the parsec level, which is much larger than the envelope scales we focus on, it is expected that $\overline{T}_{\rm env}$ should be comparable to or higher than $T_{\rm c}$, depending on whether the bolometric luminosity is low or high. One should note that $T_{\rm c}$ is not available for HOPS-354, as it is outside the region covered by Lombardi et al. (2014). However, the comparison between $\overline{T}_{\rm env}$ and $T_{\rm c}$ (see the right panel of Figure 1) indicates that $\overline{T}_{\rm env}$ is within a reasonable range for all protostars. Therefore, we use $\overline{T}_{\rm env}$ to estimate the envelope physical properties for HOPS-354.
@@
Table LABEL:Tab:parameters summarizes the derived parameters for each protostar, including information from the literature. Column 2 to 8 of Table LABEL:Tab:parameters provide the following data: bolometric luminosity $L_{\rm bol}$, total flux density $S_{\nu}$, mean envelope dust temperature $\overline{T}_{\rm env}$, cloud dust temperature $T_{\rm c}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, and gas number density $n_{\rm gas}$. Figure 2(A–F) illustrates the distributions of these parameters. The median and mean values of the sample are as follows: $L_{\rm bol}$: 6.57 and 63.87 L, $S_{\nu}$: 0.51 and 0.71 Jy, $\overline{T}_{\rm env}$: 24.2 and 29.9 K, $M_{\rm env}$: 0.30 and 0.40 M, $N_{\rm gas}$: 1.5×$\times 10^{23}$ and 2.0×$\times 10^{23}$ cm-2, and $n_{\rm gas}$: 6.1×$\times 10^{6}$ and 8.2×$\times 10^{6}$ cm-3.
Figure 2: The stacked histogram of physical envelope parameters for protostars with different B-field structure. Panel A to H: bolometric luminosity $L_{\rm bol}$, mean dust envelope temperature $\overline{T}_{\rm d}$, total flux density $S_{\nu}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, gas number density $n_{\rm gas}$, non-thermal velocity dispersion $\sigma_{\rm nth}$ and angle dispersion of B-field $\delta\phi$, respectively. Dashed and solid lines indicate the mean and median values of the distribution.
@@ 3.2 Line Width
where $\sigma_{\rm v}$ is the velocity dispersion along the line of sight, $\sigma_{\rm obs}=\Delta V_{\rm FWHM}/(2\sqrt{2\rm ln2})$, and $\Delta v_{\rm r}$ is the spectral resolution of the observations $\sim 0.87\ \rm km\ s^{-1}$. The non-thermal component of the velocity dispersion can be expressed as:
@@
where $\sigma_{\rm nth}$ and $\sigma_{\rm th}$ are the non-thermal and the thermal velocity dispersion, respectively. $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm obs}$ is the mass of the observed molecule (for C17O, $m_{\rm obs}=29\ m_{\rm H}$). $T_{\rm kin}$ is the kinetic temperature of the gas, which we assume to be equal to $\overline{T}_{\rm env}$. Due to the limited velocity resolution, only a small subset of sources had their velocity dispersions resolved in our observation. Kim et al. (2020) reported velocity dispersion values of 0.68, 0.72, 0.65, and 0.70 km s-1, which were traced by the N2H+ line on the same position as HOPS-11, HOPS-164, HOPS-399, and HOPS-403, respectively. These values are comparable to our derived dispersions (see Table LABEL:Tab:parameters), with a difference of $\sim$10%, although their studies focused on larger scales. We utilize the derived velocity dispersions for the following analysis, as our objective is to identify correlations between pairs of parameters, including the non-thermal velocity dispersion. The non-thermal velocity dispersions are listed in column 9 of Table LABEL:Tab:parameters. As shown in Figure 2(G), the median and mean value of velocity dispersion are 0.77 and 0.82 km s-1, respectively.
@@ 3.3 Angle Dispersion of B-field
The conventional approach for estimating the angle dispersion of B-fields ($\delta\phi$) is through the measurement of its standard deviation. However, this approach is limited by the 180 ambiguity in polarization angle, which complicates the accurate estimation of angle dispersion. To address this challenge, further methods have been proposed to quantify $\delta\phi$. Heitsch et al. (2001) proposed a variation in which the restriction for large angles was removed. Later Hildebrand et al. (2009) and Houde et al. (2009) applied a more sophisticated technique which uses an analytical derivation of the dispersion in polarization angles within a turbulent molecular cloud. In a recent study, Cortés et al. (2024) proposed an alternative approach to derive the angle dispersion by assuming that the local turbulence perturbing the field follows a Gaussian distribution, and they suggested that 16 Nyquist-sampling segments of B-field are required for the Gaussian distribution under the assumption of 5° for the margin of error and 10° for the mean standard deviation of polarization angles. In this study, we follow the typical method to estimate the angle dispersion $\delta\phi$ by calculating the standard deviation of the distribution of B-field segments. In order to reduce the ambiguity associated with the 180° orientation of the B-field, the distribution of the position angles is shifted when the angles are centred near 0° or 180°. The standard deviation of the shifted distribution is regarded as $\delta\phi$. To avoid the contamination of self-scattering (Huang et al., 2024), the polarization segments within the innermost 400 au region ($\sim 1\arcsec$, corresponding to the beam size) have been masked. We also mask the polarization data for regions beyond a radius of 1200 au, consistent with the scales used to derive the physical parameters in Section 3.1. Protostellar envelopes with less than 16 Nyquist-sampling segments of B-field within 1200 au regions have large uncertainties ($>5^{\circ}$) according to Cortés et al. (2024), and were therefore excluded from our analysis of angle dispersion. Among the remaining 18 protostellar envelopes with sufficient B-field segments, 5 exhibit B-field morphology that is standard-hourglass, 7 are rotated-hourglass, 4 are spiral, and 2 are complex. The angle dispersions $\delta\phi$ for these envelopes are listed in column 10 of Table LABEL:Tab:parameters, and Figure 2(H) shows the distribution of $\delta\phi$ for each type of B-field morphology. Notably, standard-hourglass envelopes tend to have lower angle dispersions. The median and mean values of $\delta\phi$ are 31.9 and 30.1, respectively.

### v2

@@ 3.1 Physical Envelope Properties
HH212M 14.00 0.52 29.1 17.2 $\pm$ 0.7 0.32 1.6E+23 6.6E+06 0.39 $\pm$ 0.15 / 1.00 /
HOPS-10 3.33 0.17 21.8 17.6 $\pm$ 0.6 0.12 6.2E+22 2.6E+06 0.49 $\pm$ 0.12 / 1.00 /
HOPS-12E 7.31 0.33 24.8 18.2 $\pm$ 1.4 0.21 1.0E+23 4.3E+06 0.51 $\pm$ 0.23 / 2.50 /
HOPS-50 4.20 0.32 22.4 23.8 $\pm$ 1.9 0.23 1.2E+23 4.8E+06 0.64 $\pm$ 0.29 / 1.00 /
HOPS-53 26.42 0.37 34.1 36.8 $\pm$ 4.0 0.15 7.6E+22 3.2E+06 0.36 $\pm$ 0.15 / 1.00 /
HOPS-60 21.93 0.55 32.6 27.5 $\pm$ 4.7 0.24 1.2E+23 5.0E+06 0.92 $\pm$ 0.30 / 1.00 /
HOPS-81 1.24 0.14 20.5 22.1 $\pm$ 1.6 0.11 5.6E+22 2.4E+06 0.56 $\pm$ 0.20 / 1.00 /
HOPS-84 49.11 0.59 39.8 21.7 $\pm$ 1.7 0.20 1.1E+23 4.2E+06 0.39 $\pm$ 0.23 / 2.00 /
HOPS-153 4.43 0.22 22.5 15.1 $\pm$ 1.1 0.16 7.8E+22 3.3E+06 0.87 $\pm$ 0.27 / 1.00 /
HOPS-203N 20.44 0.42 32.0 16.0 $\pm$ 1.1 0.18 9.0E+22 3.8E+06 0.50 $\pm$ 0.19 / 2.50 /
HOPS-203S 20.44 0.38 32.0 16.0 $\pm$ 1.1 0.16 8.1E+22 3.4E+06 0.43 $\pm$ 0.20 / 2.00 /
HOPS-247 3.09 0.37 21.6 13.4 $\pm$ 0.7 0.34 1.7E+23 7.1E+06 0.74 $\pm$ 0.30 / 1.00 /
HOPS-340 1.85 0.21 20.8 16.0 $\pm$ 0.8 0.20 1.0E+23 4.2E+06 0.36 $\pm$ 0.22 / 1.25 /
HOPS-354 6.57 0.34 24.2 16.2 $\pm$ 0.1 0.18 9.1E+22 3.8E+06 0.54 $\pm$ 0.23 / 1.00 /
HOPS-358 24.96 0.44 33.7 14.7 $\pm$ 1.0 0.22 1.1E+23 4.6E+06 1.34 $\pm$ 0.56 / 1.00 /
HOPS-383 7.83 0.22 25.2 23.4 $\pm$ 3.0 0.14 6.9E+22 2.9E+06 0.70 $\pm$ 0.34 / 1.00 /
Column 2 to 12 present the bolometric luminosity ($L_{\rm{bol}}$), total continuum flux density at 345 GHz ($S_{\nu}$), averaged envelope dust temperature ($\overline{T}_{\rm env}$), cloud dust temperature ($T_{\rm c}$), envelope gas mass ($M_{\rm env}$), gas column density ($N_{\rm gas}$), gas number density ($n_{\rm gas}$), non-thermal velocity dispersion ($\sigma_{\rm nth}$), B-field angle dispersion ($\delta\phi$), fragmentation level ($N_{F}$), and number of Nyquist-sampling B-field segments ($N_{B}$), respectively. “Standard”, “Rotated”, and “Spiral” indicate protostars with different types of B-field structure of standard hourglass, rotated hourglass, and spiral configuration, respectively. “Others” means the B-field pattern is complex or insufficient to identify its morphology. “Scattering” represents protostars with compact polarized emission, which is thought to be produced by self-scattering (Huang et al., 2024).$\dagger$\$$\dagger$\$footnotemark: Typical uncertainties for the different parameters are: 10% for $S_{\nu}$; 20–30% for $\overline{T}_{\rm env}$; 30–40% for $M_{\rm env}$, $N_{\rm gas}$, and $n_{\rm gas}$.
@@
In the case of a dusty cloud heated by internal stellar radiation, the radial temperature profile can be expressed as a power law: $T_{\rm dust}\propto r^{-2/(4+\beta)}$ (Kenyon et al., 1993), where $\beta$ is the dust opacity spectral index. Girart et al. (2009) adopted a value of $\beta=1.5$ when studying low-mass protostars at 1000 au scales. Kwon et al. (2009) suggested that $\beta$ should be around or less than 1 for Class 0 protostars. Tobin et al. (2020) fitted a temperature power-law index of -0.46 for hundreds-au-size disks within the Orion star-forming protostellar system, corresponding to $\beta\sim 0.35$, which is scaled by luminosity. However, this profile shows a slight decline in the outer 100 au regions due to backheating from the surrounding envelope (Tobin et al., 2020), it is subsequently recommended to refine the index to -0.50 (Sheehan et al., 2022). For this study, we use the temperature model proposed by Tobin et al. (2020) for Orion star-forming regions, with $\beta$ set to 1 (corresponding to the index of $\sim-0.40$) for BOPS protostellar envelopes (most of them are in Class 0 stage). Therefore, the dust temperature $T_{\rm dust}$ can be expressed as:
@@
where $T_{0}=43$ K is the average temperature for a $\sim$ 1 $L_{\odot}$ protostar at a radius of $\sim$50 au (Whitney et al., 2003; Tobin et al., 2013, 2020), $r$ is the distance from the protostar, and $L_{\rm bol}$ is the bolometric luminosity taken from Furlan et al. (2017). Given the clear detection of C17O emission in our sample, it is inferred that the dust temperature cannot fall below 20 K, as CO would otherwise be frozen onto the dust grains, making it undetectable (e.g., Jørgensen et al., 2015). The Equation 2 may yield very low dust temperatures (below 20 K) at the edge of the selected regions, especially when the bolometric luminosity $L_{\rm bol}$ is very low. In such cases, we use 20 K as the minimum dust temperature for pixels where the calculated temperature falls below this value. Excluding the extremely high temperature in the central pixel, we calculate the average envelope dust temperature $\overline{T}_{\rm env}$ within inner regions with a radius of $R=1200$ au.
Lombardi et al. (2014) used Planck and Herschel observations of Orion to construct and fit SEDs, and further to obtain a map of dust temperature. As shown in Figure 1, comparing the cloud dust temperature $T_{\rm c}$ (Lombardi et al., 2014) with the envelope dust temperature $\overline{T}_{\rm env}$ (this work, derived using our method described above), we find $\overline{T}_{\rm env}$ in most BOPS protostars is larger than $T_{\rm c}$, while 10 protostars have values of $\overline{T}_{\rm env}$ comparable to the dust temperature of the surrounding molecular cloud ($|\overline{T}_{\rm env}-T_{\rm c}|\lesssim 3$, the value of 3 is the half of the mean error for $|\overline{T}_{\rm env}-T_{\rm c}|$, see the notemark of Table LABEL:Tab:paras or Section 4 for error discussions). Given that the scales of molecular clouds in question are on the parsec level, which is much larger than the envelope scales we focus on, it is expected that $\overline{T}_{\rm env}$ should be comparable to or higher than $T_{\rm c}$. However, HOPS-409 shows that the value of $\overline{T}_{\rm env}$ is lower than that of $T_{\rm c}$, probably due to the cloud providing shielding from the interstellar radiation field. We note that 6 protostars lack information of $\overline{T}_{\rm env}$, in this case, $T_{\rm c}$ is employed to estimate the physical properties of the envelope instead.
@@
Table LABEL:Tab:paras summarizes the derived parameters for each protostar, including information from the literature. Column 2 to 8 of Table LABEL:Tab:paras provide the following data: bolometric luminosity $L_{\rm bol}$, total flux density $S_{\nu}$, mean envelope dust temperature $\overline{T}_{\rm env}$, cloud dust temperature $T_{\rm c}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, and gas number density $n_{\rm gas}$. Figure 2(A–B) illustrate the distributions of bolometric luminosity and envelope dust temperature for 55 protostars (the remaining 6 lack luminosity information), while Figure 2(C–F) present the distributions for parameters of flux intensity, envelope mass, gas column density and number density for 61 targets.
Figure 2: The stacked histogram of physical envelope parameters for protostars with different B-field structure. Panel A to H: bolometric luminosity $L_{\rm bol}$, mean dust envelope temperature $\overline{T}_{\rm env}$, total flux density $S_{\nu}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, gas number density $n_{\rm gas}$, non-thermal velocity dispersion $\sigma_{\rm nth}$ and angle dispersion of B-field $\delta\phi$, respectively. Dashed lines indicate the mean values of the distribution for protostars with different types of B-field structure.
@@ 3.2 Line Width
where $\sigma_{\rm v}$ is the velocity dispersion along the line of sight, and $\Delta v_{\rm r}$ is the spectral resolution of the observations $\sim 0.87\ \rm km\ s^{-1}$. The observed velocity dispersion $\sigma_{\rm obs}=\Delta V_{\rm FWHM}/(2\sqrt{2\rm ln2})$, where $\Delta V_{\rm FWHM}$ is estimated using the moment one map of C17O within the same regions as that for the estimation of the physical parameters. In light of the potential impact of the gravity and outflows on observed dispersion, we exculded the innermost regions with a radius of 200 au around the dust peak, and the region where the CO outflow emission is larger than 5-$\sigma$. Then the non-thermal component of the velocity dispersion can be expressed as:
@@
where $\sigma_{\rm nth}$ and $\sigma_{\rm th}$ are the non-thermal and the thermal velocity dispersion, respectively. $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm obs}$ is the mass of the observed molecule (for C17O, $m_{\rm obs}=29\ m_{\rm H}$). $T_{\rm kin}$ is the kinetic temperature of the gas, which we assume to be equal to $\overline{T}_{\rm env}$. The non-thermal velocity dispersions are listed in column 9 of Table LABEL:Tab:paras. Figure 2(G) shows the distribution of $\sigma_{\rm nth}$ for 60 protostars (HOPS-408 has no detection of velocity dispersion within the selected region).
@@ 3.3 Angle Dispersion of B-field
The conventional approach for estimating the angle dispersion of B-fields ($\delta\phi$) is by measuring its standard deviation. However, this approach is limited by the 180 ambiguity in polarization angle, which complicates the accurate estimation of the angle dispersion. To address this challenge, further methods have been proposed to quantify $\delta\phi$. Heitsch et al. (2001) proposed a variation in which the restriction for large angles was removed. Later Hildebrand et al. (2009) and Houde et al. (2009) applied a more sophisticated technique which uses an analytical derivation of the dispersion in polarization angles within a turbulent molecular cloud. In a recent study, Cortés et al. (2024) proposed an alternative approach to derive the angle dispersion by assuming that the local turbulence perturbing the field follows a Gaussian distribution, and they suggested that 16 Nyquist-sampling segments of B-field are required for the Gaussian distribution under the assumption of 5° for the margin of error and 10° for the mean standard deviation of polarization angles. In this paper, we follow the typical method to estimate the angle dispersion $\delta\phi$ by calculating the standard deviation of the distribution of B-field segments. In order to reduce the ambiguity associated with the 180° orientation of the B-field, the distribution of the position angles is shifted when the angles are centred near 0° or 180°. The standard deviation of the shifted distribution is regarded as $\delta\phi$. To avoid the contamination of self-scattering (Huang et al., 2024), the polarization segments within the innermost 400 au region ($\sim 1\arcsec$, corresponding to the beam size) have been masked. We also mask the polarization data for regions beyond a radius of 1200 au, consistent with the scales used to derive the physical parameters in Section 3.1. Protostellar envelopes with less than 16 Nyquist-sampling segments of B-field within 1200 au regions have large uncertainties ($>5^{\circ}$) according to Cortés et al. (2024), therefore we only estimate the angle dispersion of B-field for 25 protostars with sufficient statastics ($>$ 16 Nyquist-sampling segments), which is listed in column 10 of Table LABEL:Tab:paras. Among these 25 protostellar envelopes, 6 exhibit B-field morphology that is standard-hourglass, 8 are rotated-hourglass, 4 are spiral, and 7 are complex or lack polarized emission to identify their B-field morphology. Figure 2(H) shows the distribution of $\delta\phi$ for each type of B-field morphology.